Zeronknight (zeronknight):

Match this equality with the order in which you would evaluate a given value for x. 18x - 12 > 24 A. Multiply then subtract B. Divide then add C. Add then divide D. Subtract then multiply

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (dayakar):

do u have any idea

Zeronknight (zeronknight):

D?

OpenStudy (aravindg):

Why do you think D? Can you show your step using D?

Zeronknight (zeronknight):

I know people would usually subtract in problems like these. 18 - 12 = 6 24 - 12 = 24

Zeronknight (zeronknight):

I don't know if multiplication is correct

Zeronknight (zeronknight):

WAIT NO IT'S ADDITION

OpenStudy (aravindg):

Don't assume. Understand the procedure! We need to separate constants and x first.

Zeronknight (zeronknight):

I keep forgetting that's a negiative 12! we need to cancel out the 12 12 + 12 = 0 24+12=36

OpenStudy (aravindg):

Right you are!

Zeronknight (zeronknight):

I just hate how the minus sign isn't next to the 12 qnq

OpenStudy (aravindg):

Now all you gotta do is remove the coeff of x.

Zeronknight (zeronknight):

Even thou the answer is C. Why would I divide in this problem?

Zeronknight (zeronknight):

To cancel out 18?

Zeronknight (zeronknight):

@AravindG

OpenStudy (briellshaw):

The answer is C i think i am working on the same thing in school right now

OpenStudy (aravindg):

Indeed my friend! :)

Zeronknight (zeronknight):

Could you please not give out direct answers? Also, I have been helped already.

OpenStudy (briellshaw):

You should have known the answer alll along

Zeronknight (zeronknight):

I am not going to argue with you. I am probably younger than you and I am acting more mature.

OpenStudy (briellshaw):

I am 11

Zeronknight (zeronknight):

Yeah I am not going ignore you. But this is openSTUDY not openANSWERS. Please do not give out any direct answers. And I am already being helped.

OpenStudy (briellshaw):

Ok what ever do you go to Agora

OpenStudy (briellshaw):

What grade are u in

Zeronknight (zeronknight):

Why do you ask?

OpenStudy (aravindg):

@Zeronknight Don't worry about him. He just got a break :)

Zeronknight (zeronknight):

@AravindG XD. Thank you. And I probably handled that situation poorly. Sorry for my part in that argument.

OpenStudy (aravindg):

It's alright! I am happy that you did tell him about 'how not to answer questions here' ;)

Zeronknight (zeronknight):

XD. And thank you for helping me, with the question!

Zeronknight (zeronknight):

I didn't report her?

Zeronknight (zeronknight):

Are you trying to start something?

Zeronknight (zeronknight):

@AravindG

Zeronknight (zeronknight):

it's not to smart to cuss at a mod ;)

Zeronknight (zeronknight):

I see. Um how exactly will you beat this reatrice?

Zeronknight (zeronknight):

@AravindG Are u there? owo

OpenStudy (aravindg):

Sorry about that @Zeronknight My pc is quite slow considering the fact that it runs on the old Pentium processor. Sorry for the delay. Necessary action has been taken.

Zeronknight (zeronknight):

XD. And it'[s fine. And alright! Thank you again!

OpenStudy (aravindg):

yw :)

OpenStudy (briellshaw):

Um wow i didn't know that my little brother was online i was in the bathroom when he was and when i came back to log on it said i was suspended i am sorry if you are confused earlier that was not me cursing and stuff it wasn't even me when i was suspended i had two accounts because sometime i forget the password to this one and now i see i cant even log on that one again i am so very sorry for my Little brothers actions